Cisplatin is found to be 65.02 % Pt, 23.63 % Cl, 9.34 % N, and 2.02 % H by mass. Determine its empirical formula.



1
Expert's answer
2021-03-24T06:49:28-0400

65.02g Pt × 1 mol Pt /195.1g = 0.3333 mol Pt

9.34g N × 1 mol N/ 14.06 g = 0.667 mol N

2.02g H × 1 mol H/ 1.01 g = 2.00 mol H

23.63g Cl ×1 mol /35.45g = 0.6666mol Cl

Pt 0.3333/0.3333 =. 1

N 0.667 / 0.3333 = 2

H 2.00/0.3333 = 6

Cl = 0.6666/0.3333 = 2

EF = P + N2 + H6 + Cl2